OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x+y= -1 x-3y= 11 Please tell me how you got your answer- thank you for helping me! :) A. (2, −3) B. (−2, 1) C. (5, −2) D. (−6, 5)

OpenStudy (greatlife44):

You can solve this by a substitution method or eliminating one of the variables

OpenStudy (greatlife44):

the easy way would be to plug in the answer choices wherever you see x and y

OpenStudy (cathyangs):

If you have trouble rearranging systems of equations, you can simply test each of the ordered pairs, using the x-value in place of x and the y-value in place of y, to see which pair holds true for both equations. If you want, I can help you work through them, though, which is probably what the person giving you this question wanted you to do.

OpenStudy (anonymous):

Yes please!

OpenStudy (greatlife44):

Take a look at this x+y = -1 y = -1-x x -3(-1-x) = 11 x +3+3x = 11 4x+3 = 11 -3 -3 4x = 8 x = 2 x + y = -1 2+y = -1 y = -1-2 y = -3 2-3(-3) 2-(-9) = 11 so x = 2 and y = -3

OpenStudy (anonymous):

Thank you so much!

OpenStudy (greatlife44):

ideally you don't want to just plug in the answer choices because you won't always be given both x and y

OpenStudy (greatlife44):

of course when you solve this yourself, you can plug in at the end.

OpenStudy (anonymous):

I'll keep that in mind, thank you :) @greatlife44